# Statistical Principles for Platform Trials

### v4

@@ 6 Conclusion
A significant advantage of a Platform trial is that multiple treatments are evaluated simultaneously against a shared control, which allows for fewer total participants, and greater allocation of participants to the active treatment within each regimen, relative to traditional designs with 1:1 randomization. However, the stability of regulatory FAR across Platform trials is of serious concern since shared controls introduce dependencies among treatment evaluations and therefore increases the variability of ERpf. As we demonstrated in our analysis, dependencies among treatment evaluations increases with the treatment/control randomization ratio. From patients perspective, a higher treatment/control randomization ratio will allow higher chance of receiving a new treatment compound. From an individual drug sponsor’s point of view, a higher treatment/control randomization ratio through participating platform trial allows more accurate estimate on the effects of new compounds. However, from the master protocol sponsor’s point of view, lower the the treatment/control randomization ratio allows increasing the sample size of the shared control, and therefore lessens the chance that an overachieving shared control sample causes all the new compounds have difficulty to demonstrate efficacy. Additionally, from the regulators’ point of view, stability of ERpF control is affected by both stability of the sample mean of the shared control as well as correlation among the pivotal statistics. A lower randomization ratio increases sample size of the shared control, which lessens the chance of an under-performing shared control sample giving all the new compounds an easy pass. A lower randomization ratio also decreases correlation among the pivotal statistics, which makes ERpF control more stable. To balance the different perspective from different stake holder, our Confident RandomSizer app offers a tool to identify optimal treatment/control randomization ratio given a fixed sample size.
@@ 7.1 Terms
  • FAR: False Approval Rate; proportion of drug compounds that lack efficacy incorrectly approved
@@ 7.2 Notation
  • $\mu_{SC}$: Shared control sample mean
@@
  • $\hat{\mu}_{SC}^{\star}$: Standardized shared control sample mean estimate
@@
  • $k$: Number of studies

### v5

@@ 6 Conclusion
The Conditionality Principle does not weaken frequentist guarantees; it rather aligns inference with the experiment that actually occurred.
  1. 1.
    Among error rates defined by Tukey (1953), Error Rate Familywise (ERFw) applies within a clinical study, while Error Rate per Family (ERpF) applies across studies.
  2. 2.
    An appropriate version of the Law of Large Numbers shows controlling ERFw within a clinical study at $5\%$ in turn controls ERpF across studies at 5-per-100, regardless of whether the studies are in Platform trials or not.
  3. 3.
    With ongoing regulatory practice controlling ERpF across studies at 5-per-100 (via controlling the ERFw of each clinical study at $5\%$, the False Approval Rate (FAR) is being controlled at 5-per-100 (even in Platform trials), allowing only five incorrect approvals per 100 drug compounds that lack efficacy.
  4. 4.
    False Approval Rate (FAR), the proportion of drug compounds lacking efficacy incorrectly approved by regulators, should not be increased with increasing number of efficacious compounds approved (according to our understanding of the U.S. Food, Drug, & Cosmetic Act and its Amendment).
  5. 5.
    Controlling the False Discovery Rate (FDR) across studies at $5\%$ would make FAR more than 5-per-100, a change of (inter)national policies more than statistical that we believe would require harmonizsation/harmonization.
@@
A significant advantage of a Platform trial is that multiple treatments are evaluated simultaneously against a shared control, which allows for fewer total participants, and greater allocation of participants to the active treatment within each regimen, relative to traditional designs with 1:1 randomization. However, the stability of regulatory FAR across Platform trials is of serious concern since shared controls introduce dependencies among treatment evaluations and therefore increases the variability of ERpf. As we demonstrated in our analysis, dependencies among treatment evaluations increases with the treatment/control randomization ratio. From patients perspective, a higher treatment/control randomization ratio will allow higher chance of receiving a new treatment compound. From an individual drug sponsor’s point of view, a higher treatment/control randomization ratio through participating platform trial allows more accurate estimate on the effects of new compounds. However, from the master protocol sponsor’s point of view, lower the treatment/control randomization ratio allows increasing the sample size of the shared control, and therefore lessens the chance that an overachieving shared control sample causes all the new compounds have difficulty to demonstrate efficacy. Additionally, from the regulators’ point of view, stability of ERpF control is affected by both stability of the sample mean of the shared control as well as correlation among the pivotal statistics. A lower randomization ratio increases sample size of the shared control, which lessens the chance of an under-performing shared control sample giving all the new compounds an easy pass. A lower randomization ratio also decreases correlation among the pivotal statistics, which makes ERpF control more stable. To balance the different perspective from different stake holder, our Confident RandomSizer app offers a tool to identify optimal treatment/control randomization ratio given a fixed sample size.
@@
7 Conflict of Interest
X.C. and E.Ou declare no competing interests. Y. L. is an employee and shareholder at Nektar Therapeutics, San Francisco, CA, USA. J.Y.S and H. T. are employees and shareholders of BeOne Medicines (San Mateo, CA, USA). B.W. is an employee at IO Biotech (Copenhagen, Denmark). J.C. H. is a professor emeritus of the Ohio State University and served as a consultant to BeOne Medicines and Eli Lilly.
@@ 8.1 Terms
  • IAR: Incorrect Approval Rate; proportion of drug compounds that lack efficacy incorrectly approved
@@ 8.2 Notation
  • $\mu_{SC}$: Shared control true sample mean
@@
  • $k$: Number of treatment arms
  • $\ell$: Number of non-efficatious treatment arms